OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by y=x^2 , y = 0, and x = 2,, about the y-axis. I get 6.2831853, but its wrong. Help me understand where I went wrong, please.

OpenStudy (anonymous):

x=sqrty intgrated[0,2] pi(y)dy = 6.2831853

OpenStudy (campbell_st):

well not quite you are rotating about the y-axis so substitute x =2 to find the corresponding y value  then its \[V=\pi \int\limits_{0}^{4} x^2 dy\] you know \[y = x^2\] so substitute and you have \[V = \pi \int\limits_{0}^{4}(y) dy\] hopefully you can get the rest from here...